# Primordial black hole formation from self-resonant preheating?

### v1

@@ I Introduction
More recently, the matter has received renewed attention, specifically in the context of so-called metric preheating in single-field inflation Martin:2019nuw ; Auclair:2020csm . These works have argued that PBH formation from metric preheating is basically unavoidable. This would be in stark contrast with the tuning that is generically required to produce them in models in which they originate from large curvature fluctuations generated during inflation. Moreover, these articles conclude that the mechanism is so efficient that those PBHs can dominate the energy density of the Universe after inflation and reheating can proceed thanks to Hawking evaporation. The PBH originating from preheating would have masses too small to account for the current dark matter – a strong motivation for current PBH studies, see Carr:2020xqk ; Green:2020jor ; Sasaki:2018dmp ; Escriva:2022duf ; Villanueva-Domingo:2021spv for recent reviews including bounds on the abundance of these objects –, as they would have evaporated before today. Nevertheless, their interest for cosmology remains, e.g. they could alter the evolution of the Universe with respect to the standard picture and even be catalysers for the generation of the dark matter Khlopov:2004tn ; Fujita:2014hha ; Lennon:2017tqq ; Hooper:2019gtx ; Baldes:2020nuv ; Bernal:2021bbv ; Masina:2021zpu . In particular, they would suggest the possibility of a much more important role for PBHs than it is currently accepted.
@@
In this work, we first revisit these claims that metric preheating leads to generic PBH production thanks to the enhanced growth of scalar fluctuations. Then we address the question of whether a generalised scenario for self-resonant formation of PBH during reheating is viable. Our analysis is structured as follows. In Section II we review the mechanism of parametric resonance during preheating and discuss the evolution of the inflation perturbations in a toy model based on a purely quadratic monomial inflationary potential, as it was considered in Martin:2019nuw ; Auclair:2020csm . In Sec. II.1 we discuss the observational limitations of this scenario as well as its inability to produce PBHs, in the light of existing literature. In Section III we focus on more realistic scenarios: On the one hand, we show that accounting for anharmonic terms is typically needed (and these terms are dominant with respect to gravitational ones) in a vast part of the parameter space of viable inflationary models, including popular benchmark ones like Starobinsky inflation Starobinsky:1980te . Thus, the simple scenario described in Sec. II is not only strictly speaking ruled out, but cannot be considered as representative of inflationary models, resulting far from generic. On the other hand, one might wonder if anharmonic terms may rescue PBH production from preheating in a modified version of the scenario. Specifically, one may ask if inflaton self-interactions may boost the formation of non-linear structures if compared to pure gravitational interaction, thus easing PBH formation. After studying the dynamics in the linear regime (Sec. III), in Section IV we follow the evolution of the perturbations in the nonlinear regime with the help of lattice simulations. We find that the parameter space where the formation of non-linear structures is significantly enhanced corresponds to the one where these objects are metastable transients or oscillons (see e.g. Bogolyubsky:1976yu ; Gleiser:1993pt ; Copeland:1995fq ; Honda:2001xg ; Saffin:2006yk ; Hindmarsh:2007jb ; Amin:2010jq ; Amin:2010dc ), rather than PBHs: in this range, we recover results previously obtained in the context of early universe oscillon physics. Our study thus implies that PBH formation from preheating dynamics is challenging and certainly non-generic. In Section V, we review our results, discuss some perspectives and conclude. Some considerations on the quantitative reliability with which the post-inflationary expansion can be approximated as matter-like are reported in the appendix.
@@ II Parametric resonance for the quadratic potential: Metric preheating
$\omega^{2}=\frac{{\rm d}^{2}V}{{\rm d}\phi^{2}}+\frac{k^{2}}{a^{2}}+\frac{2}{M_{\rm Pl}^{2}}\frac{{\rm d}V}{{\rm d}\phi}\frac{\dot{\phi}}{H}+\frac{3\dot{\phi}^{2}}{M_{\rm Pl}^{2}}-\frac{\dot{\phi}^{4}}{2H^{2}M_{\rm Pl}^{4}}+\frac{3}{4M_{\rm Pl}^{2}}P\,,$ (9)
where $P$ denotes the background pressure $P=\dot{\phi}/2-V$.
A quadratic potential, $V=\frac{1}{2}m^{2}\phi^{2}$, is sometimes used to approximate the dynamics of the inflaton background during reheating. This approximation can be useful provided that the oscillations of the inflaton condensate about the minimum of the potential, once inflation has ended, are small enough. Assuming such a potential, the condition $\epsilon=1$ gives the value of $H$ at the end of inflation, $H_{\rm end}=m\,\phi_{\rm end}/(2M_{\rm P})$. Then, approximating $\epsilon$ by $\epsilon_{V}$, we get $\phi_{\rm end}\simeq\sqrt{2}M_{\rm P}$, which leads to $H_{\rm end}\simeq m/{\sqrt{2}}$. If we denote with $t_{\rm end}$ the (cosmic) time at which inflation ends, the following asymptotic solution of the homogeneous inflaton field equation $\ddot{\phi}+3H\dot{\phi}+{\rm d}V/{\rm d}\phi=0$ holds for times $t\gg t_{\rm end}$ so that $H\ll m$:
@@
Figure 1: Post-inflationary time evolution of some representative terms in the parenthesis in Eq. (9), normalised to $m^{2}$, for a quadratic potential with $m=1.56\times 10^{13}$eV. We have imposed $N_{*}=55$ and a normalization of the power spectrum of curvature perturbations as explained in Sec.III.1. The numerical results are reported in red and blue for the third and fourth term in Eq. (9), respectively. The dashed lines indicate the analytical result obtained with eq. (10).
Within the regime of validity of the approximate solution eq. (10), the last three terms on the RHS of eq. (9) redshift as $a^{-3}$ and can be neglected at sufficiently late times because the third one scales as $a^{-3/2}$, see also Jedamzik:2010dq . In Figure 1 we verify this statement numerically for a reasonable choice of parameters by comparing the third term (red) with the fourth one (blue), normalised to $m^{2}$ (second derivative of the potential). The colored lines show the numerical solutions, while the dashed lines correspond to the analytical approximation of eq. (10). A few e-folds past the time at which $\epsilon=1$ (when $\Delta N=0$) are shown; this roughly corresponds to $\phi$ evolving from $\sim M_{\rm P}$ to $\sim 10^{-4}\,M_{\rm P}$.
@@
$A(k)=1+\frac{k^{2}}{m^{2}a^{2}}\quad\text{and}\quad q=\frac{\phi_{\rm end}}{M_{\rm Pl}}\left(\frac{a_{\rm end}}{a}\right)^{3/2}\,.$ (12)
@@
For $A(k)$ and $q$ constants, Eq. (11) is familiar in inflationary preheating theory Kofman:1994rk ; Kofman:1997yn ; it admits solutions of the form Amin:2014eta ; lozanov2019lectures
@@
with $\mathcal{P}_{k\pm}$ periodic functions of $s$ with period equal to the one of the oscillating condensate, $T=2\pi/m$; $\mu_{k}$ is the complex-valued Floquet exponent and depends on both $A$ and $q$. Regions of the parameter space $\{A,q\}$ where $\Re(\mu_{k})\neq 0$, usually referred to as bands, correspond to an exponential increase with time of one of the solutions. This phenomenon is dubbed parametric resonance and describes a fast growth of perturbations with time at the expense of the the inflaton condensate oscillating at the bottom of the potential.
@@
the narrow resonance condition [^3] In fact, the narrow resonance condition $q\ll 1$ is not fulfilled soon after the end of inflation, when $q_{\rm end}\simeq\sqrt{2}$. The reliability of the adiabatic approximation in the linear theory around the unaffected background homogeneous field is then unclear, being controlled by the time- and mode-varying value of $\Re(\mu_{k})m/H$. The applicability of the analytical theory assuming $q\ll 1$ everywhere in articles such as Jedamzik:2010dq thus relies fully on the numerical solution. yields $\Re(\mu_{k})=|q|/2$, the growth of $\tilde{v}_{k}$ can be heuristically obtained by the “adiabatic” approximation $\tilde{v}_{k}\propto\exp({\rm d}s\int\mu_{k})\propto a^{3/2}$ Finelli:1998bu ; Jedamzik:2010dq ; Hertzberg:2014jza , hence one derives that the density contrast grows as $a$ in the perturbative regime Nambu:1996gf ; Finelli:1998bu ; Jedamzik:2010dq , recovering the behaviour of dust in Newtonian linear perturbation theory, as it was remarked almost forty years ago 1985MNRAS.215..575K .
@@ II.1 Is the metric preheating instability leading to PBH formation?
There exists consensus in the literature about the previous result, at least within the regime of validity of the approximations made (more on their realism/plausibility later). However, at times it has been the basis for much bolder statements concerning the fate of these growing density contrast modes, which have been argued to lead quite generically to PBH formation. [^4] In order to avoid confusion, let us clarify that the mechanism of interest here is not the same considered e.g. in Green:2000he ; Bassett:2000ha , which instead concerns two-scalar field models, and the modes considered for collapse are Hubble-sized. Lattice studies have shown, by the way, that even in this case PBH production is typically not relevant Suyama:2004mz ; see however Torres-Lomas:2014bua for considerations on sub-Hubble modes. It also differs from tachyonic preheating, i.e. where the instability is driven by the sign of ${\rm d}^{2}V/{\rm d}\phi^{2}$ dynamically turning negative, which was again found to be inconsequential for PBH production in models consistent with CMB data Suyama:2006sr . Finally, “axion” inflation coupled to gauge fields has also been revealed less prone to PBH production Caravano:2022epk than initially thought Linde:2012bt ; Garcia-Bellido:2016dkw .
@@
This was presented as a possibility already in Jedamzik:2010dq , but as far as we can infer, it was not based on an analysis of the non-linear evolution of the modes. We remind the reader that a well-accepted criterion for BH formation is that the hoop conjecture should be verified (see Ref. Misner:1973prb , pp. 867-868): the entirety of the object’s mass must be compressed to the point that it can fit within a sphere whose radius is equal to that object’s Schwarzschild radius. This is loosely equivalent to requiring that the Newtonian potential $|\Phi|\sim GM/R$ associated to the mass $M$ concentrated in a volume of radius $R$ attains $|\Phi|\sim 0.5$ (see e.g. Lyth:2005ze ). [^5] Some literature such as Cotner:2018vug ; del-Corral:2023apl has sometimes estimated PBH formation at sub-horizon scales relying on different criteria, which do not appear generically reliable. Notice that within a purely linear analysis, as the one in Jedamzik:2010dq , there is really no indication that the growth of density perturbations may lead to BH formation, since the curvature perturbation $\mathcal{R}$ (reducing in modulus to the gravitational potential $\Phi$ on sub-Hubble scales, see footnote 2) stays constant: Put otherwise, extrapolating the linear behaviour into the non-linear regime (a procedure which is not anyway to be trusted Parry:1998pn ) is not indicative of PBH formation.
@@
In the appendices of Martin:2019nuw , the allegedly rather generic formation of PBHs was supported by solving the non-linear evolution equation of a spherically symmetric fluctuation of the energy density of the field, evolving in the homogeneous background of the scalar field itself. Unfortunately, this argument is inconclusive, since it does not capture the relevant physics of the situation that it hopes to describe. It would suffice to notice that, within the same approximation used there, any fluctuation in a dust fluid would end in a BH during matter domination, which is not the case. As noted already in the pioneering studies on the subject Khlopov:1980mg ; Polnarev:1985btg ; 1985MNRAS.215..575K , if and only if the fluctuations are homogeneous and spherically symmetric to a high-degree, the studied solutions can be taken to represent reality. A number of articles have since been devoted to clarify the parametric dependence of the probability of the formation of (Hubble-sized) PBHs on the root-mean-square relative fluctuation of density on a mass scale $M$, $\sigma(M)$, when asphericity and angular momentum are taken into account Kuhnel:2016exn ; Harada:2016mhb ; Harada:2017fjm ; Carr:2018nkm ; Harada:2022xjp . In general, collapse is prevented by a number of effects: The isotropic approximation breaks down, gradients are important, non-Gaussian density and velocity effects are present, etc.
@@
In fact, the analogous problem in cold dark matter cosmology is well-known, and the dynamics (unless the power spectrum is extremely enhanced) leads to halo formation supported by velocity dispersion, rather than to BHs. A qualitatively analogous outcome has actually been recovered in numerical simulations of early structure formation, neglecting GR effects, in a universe subject to a quadratic inflaton potential, if the post-inflationary metastable state is assumed to survive long enough Musoke:2019ima . No sign of large gravitational potentials, which would cast doubts on resorting to a Newtonian approximation, is found there. Not surprisingly, the corresponding halos have a profile resembling an NFW one Eggemeier:2020zeg . Of course, since dealing with a bosonic degree of freedom, a solitonic condensate may form at the center of halos. Even in the rather idealised case where no self-interaction or other interactions are relevant, for these objects to collapse into BHs, the matter-dominated phase must last a very long time, equivalent to $\gtrsim 30\,$ e-folds Eggemeier:2021smj . Even neglecting the physical realism of these cases, definitely the PBH outcome is far from generic. [^6] Note that, as far as BH formation is concerned, the free bosonic case these simulations focus on is optimal, with self-interactions or interactions with other fields making formation of BH more challenging, see e.g. Garani:2021gvc .
@@
Are general relativistic effects relevant to alter these conclusions? Existing studies do not support the idea that they would be responsible for BH formation: For instance, the parametric family of potentials studied in Kou:2019bbc leads to BH formation only in a subset of cases when their parameter $\alpha<0.2$ (with $\alpha=1$ corresponding to the quadratic potential, see Fig. 8 in Kou:2019bbc ). The authors of Muia:2019coe also clearly state that the fate of non-linear structures forming in a quadratic inflaton potential is to disperse, not to end up forming BHs. While this situation appears grim enough for PBH production, we note that these simulations start from already large density contrasts as initial conditions, whose dynamical plausibility in the post-inflationary era needs to be independently checked. We will return to this point in Sec. IV.

### v3

@@ I Introduction
More recently, the matter has received renewed attention, specifically in the context of so-called metric preheating in single-field inflation. Recent works have argued that PBH formation from metric preheating may be possible Auclair:2020csm or even unavoidable Martin:2019nuw under mild and rather generic conditions. This would be in stark contrast with the tuning that is generically required to produce them in models in which they originate from large curvature fluctuations generated during inflation. Moreover, these articles conclude that the mechanism is so efficient that those PBHs can dominate the energy density of the Universe after inflation and reheating can proceed thanks to Hawking evaporation. The PBH originating from preheating would have masses too small to account for the current dark matter – a strong motivation for current PBH studies, see Carr:2020xqk ; Green:2020jor ; Sasaki:2018dmp ; Escriva:2022duf ; Villanueva-Domingo:2021spv for recent reviews including bounds on the abundance of these objects –, as they would have evaporated before today. Nevertheless, their interest for cosmology remains, e.g. they could alter the evolution of the Universe with respect to the standard picture and even be catalysers for the generation of the dark matter Khlopov:2004tn ; Fujita:2014hha ; Lennon:2017tqq ; Hooper:2019gtx ; Baldes:2020nuv ; Bernal:2021bbv ; Masina:2021zpu . In particular, they would suggest the possibility of a much more important role for PBHs than it is currently accepted.
In this work, we first revisit these claims that metric preheating leads to generic PBH production thanks to the enhanced growth of scalar fluctuations. Then we address the question of whether a generalised scenario for self-resonant formation of PBH during reheating is viable. Our study implies that PBH formation from preheating dynamics is challenging and certainly non-generic. For busy readers that are familiar with the topic, we find it convenient to summarize now briefly the essence of the logic that we follow to arrive to our conclusion. We first show that the mechanism discussed in Jedamzik:2010dq ; Martin:2019nuw ; Auclair:2020csm purportedly leading to PBH formation is nothing but the linear growth of subhorizon matter fluctuations during an epoch of matter domination. Such fluctuations grow in time proportional to the scale factor of the Universe, while the curvature fluctuations remain constant. In order to draw a reliable conclusion about PBH formation in such a situation, the late-time extrapolation of the linear result is insufficient, and the existing treatments resorting to too idealized conditions, such as spherical symmetry and absence of feedback, are unreliable. Yet, the linear conservation of the curvature fluctuation already indicates that the gravitational potential is unlikely to reach the necessary values for collapse into PBHs, unless this outcome is generically rescued by non-linear effects. A shortcoming of the previous studies Martin:2019nuw ; Auclair:2020csm is that a purely quadratic potential was considered. Such a potential is strictly speaking incompatible with the CMB, hence it can only be considered as a limiting case embedded in a more complex potential. It turns out that the realistic viable choices of the potential are relevant to describe the last stages of the evolution of the inflaton and might also rescue the mechanism we are interested in, since (as we show) anharmonic effects are important in its dynamics. One may hope that these effects may enhance the growth of fluctuations during preheating, possibly favoring PBH formation, but it turns out that this is not the case. We see this by doing (non-linear) lattice computations of the evolution of the fluctutions for potentials that are known to provide a good fit to the CMB. In agreement with existing literature on the formation of non-perturbative objects in the post-inflationary epoch, the growth of the inflaton perturbations is quenched by their own backreaction, making PBH formation non-generic at best. Only the formation of fuzzier objects (oscillons) may be expected. A final hope for PBHs may be searched within the context of an analysis with full General Relativity, that would extend our lattice study. However, studies in that direction have found that PBHs could only be hoped to form under very particular initial conditions Aurrekoetxea:2023jwd , very different from the envisaged growth of generic small inflationary seeds subject to self-resonant enhancement.
@@
Our analysis (following the logic just described) is structured as follows. In Section II we review the mechanism of parametric resonance during preheating and discuss the evolution of the inflation perturbations in a toy model based on a purely quadratic monomial inflationary potential, as it was considered in Martin:2019nuw ; Auclair:2020csm . In Sec. II.1 we discuss the observational limitations of this scenario as well as its inability to produce PBHs, in the light of existing literature. In Section III we focus on more realistic scenarios: On the one hand, we show that accounting for anharmonic terms is typically needed (and these terms are dominant with respect to gravitational ones) in a vast part of the parameter space of viable inflationary models, including popular benchmark ones like Starobinsky inflation Starobinsky:1980te . Thus, the simple scenario described in Sec. II is not only strictly speaking ruled out, but cannot be considered as representative of inflationary models, resulting far from generic. On the other hand, one might wonder if anharmonic terms may rescue PBH production from preheating in a modified version of the scenario. Specifically, one may ask if inflaton self-interactions may boost the formation of non-linear structures if compared to pure gravitational interaction, thus easing PBH formation. After studying the dynamics in the linear regime (Sec. III), in Section IV we follow the evolution of the perturbations in the nonlinear regime with the help of lattice simulations. As we already mentioned, we find that the parameter space where the formation of non-linear structures is significantly enhanced corresponds to the one where these objects are metastable transients or oscillons (see e.g. Bogolyubsky:1976yu ; Gleiser:1993pt ; Copeland:1995fq ; Honda:2001xg ; Saffin:2006yk ; Hindmarsh:2007jb ; Amin:2010jq ; Amin:2010dc ), rather than PBHs: in this range, we recover results previously obtained in the context of early universe oscillon physics. In Section V, we review our results and present our conclusions. Some considerations on the quantitative reliability with which the post-inflationary expansion can be approximated as matter-like are reported in the appendix.
@@ II Parametric resonance for the quadratic potential: Metric preheating
$\omega^{2}=\frac{{\rm d}^{2}V}{{\rm d}\phi^{2}}+\frac{k^{2}}{a^{2}}+\frac{2}{M_{\rm P}^{2}}\frac{{\rm d}V}{{\rm d}\phi}\frac{\dot{\phi}}{H}+\frac{3\dot{\phi}^{2}}{M_{\rm P}^{2}}-\frac{\dot{\phi}^{4}}{2H^{2}M_{\rm P}^{4}}+\frac{3}{4M_{\rm P}^{2}}P\,,$ (9)
where $P=\dot{\phi}^{2}/2-V$ denotes the background pressure and we stress that the last three terms in (9) originate from the fluctuations of the metric.
A quadratic potential, $V=\frac{1}{2}m^{2}\phi^{2}$, is often used to approximate the dynamics of the inflaton background during reheating. This approximation can be useful provided that the oscillations of the inflaton condensate about the minimum of the potential, once inflation has ended, are small enough. Assuming such a potential, the condition $\epsilon=1$ gives the value of $H$ at the end of inflation, $H_{\rm end}=m\,\phi_{\rm end}/(2M_{\rm P})$. Then, approximating $\epsilon$ by $\epsilon_{V}$, we get $\phi_{\rm end}\simeq\sqrt{2}M_{\rm P}$, which leads to $H_{\rm end}\simeq m/{\sqrt{2}}$. If we denote with $t_{\rm end}$ the (cosmic) time at which inflation ends, the following asymptotic solution of the homogeneous inflaton field equation $\ddot{\phi}+3H\dot{\phi}+{\rm d}V/{\rm d}\phi=0$ holds for times $t\gg t_{\rm end}$ so that $H\ll m$:
@@
Figure 1: Post-inflationary time evolution of some representative terms in the parenthesis in Eq. (9), normalised to $m^{2}$, for a quadratic potential with $m=1.56\times 10^{13}$eV. We have imposed $N_{*}=55$ and a normalization of the power spectrum of curvature perturbations as explained in Sec.III.1. The numerical results are reported in red and green for the third and fourth term in Eq. (9), respectively. The dashed lines indicate the analytical result obtained with eq. (10), where we numerically solved for the scale factor.
Within the regime of validity of the approximate solution eq. (10), the last three terms on the RHS of eq. (9) redshift as $a^{-3}$ and can be neglected at sufficiently late times because the third one scales as $a^{-3/2}$, see also Jedamzik:2010dq . In Figure 1 we verify this statement numerically for a reasonable choice of parameters by comparing the third term (red) with the fourth one (green), normalised to $m^{2}$ (second derivative of the potential). The colored lines show the full numerical solutions, while the dashed lines have been obtained by solving numerically the evolution of the scale factor and plugging this quantity in the analytical approximation of eq. (10). A few e-folds past the time at which $\epsilon=1$ (when $\Delta N=0$) are shown; this roughly corresponds to $\phi$ evolving from $\sim M_{\rm P}$ to $\sim 10^{-4}\,M_{\rm P}$.
@@
$A(k)=1+\frac{k^{2}}{m^{2}a^{2}}\quad\text{and}\quad q=\frac{\phi_{\rm end}}{M_{\rm P}}\left(\frac{a_{\rm end}}{a}\right)^{3/2}\,.$ (12)
@@
For $A(k)$ and $q$ constant, Eq. (11) is familiar in inflationary preheating theory Kofman:1994rk ; Kofman:1997yn ; it admits solutions of the form Amin:2014eta ; lozanov2019lectures
@@
with $\mathcal{P}_{k\pm}$ periodic functions of $s$ with period equal to the one of the oscillating condensate, $T=2\pi/m$; $\mu_{k}$ is the complex-valued Floquet exponent and depends on both $A$ and $q$. Regions of the parameter space $\{A,q\}$ where $\Re(\mu_{k})\neq 0$, usually referred to as bands, correspond to an exponential increase with time of one of the solutions. This phenomenon is dubbed parametric resonance and describes a fast growth of perturbations with time at the expense of the inflaton condensate oscillating at the bottom of the potential.
@@
the narrow resonance condition [^3] In fact, the narrow resonance condition $q\ll 1$ is not fulfilled soon after the end of inflation, when $q_{\rm end}\simeq\sqrt{2}$. The reliability of the adiabatic approximation in the linear theory around the unaffected background homogeneous field is then unclear, being controlled by the time- and mode-varying value of $\Re(\mu_{k})m/H$. The applicability of the analytical theory assuming $q\ll 1$ everywhere in articles such as Jedamzik:2010dq thus relies fully on the numerical solution. yields $\Re(\mu_{k})=|q|/2$; the growth of $\tilde{v}_{k}$ can be heuristically obtained by the “adiabatic” approximation $\tilde{v}_{k}\propto\exp({\rm d}s\int\mu_{k})\propto a^{3/2}$ Finelli:1998bu ; Jedamzik:2010dq ; Hertzberg:2014jza , hence one derives that the density contrast grows as $a$ in the perturbative regime Nambu:1996gf ; Finelli:1998bu ; Jedamzik:2010dq , recovering the behaviour of dust in Newtonian linear perturbation theory, as it was remarked almost forty years ago 1985MNRAS.215..575K .
@@ II.1 Is the metric preheating instability leading to PBH formation?
There exists consensus in the literature about the previous results, at least within the regime of validity of the approximations made (more on their realism/plausibility later). However, at times it has been the basis for much bolder statements concerning the fate of these growing density contrast modes, which have been argued to lead quite generically to PBH formation. [^4] In order to avoid confusion, let us clarify that the mechanism of interest here is not the same considered e.g. in Green:2000he ; Bassett:2000ha , which instead concerns two-scalar field models, and the modes considered for collapse are Hubble-sized. Lattice studies have shown, by the way, that even in this case PBH production is typically not relevant Suyama:2004mz ; see however Torres-Lomas:2014bua for considerations on sub-Hubble modes. It also differs from tachyonic preheating, i.e. where the instability is driven by the sign of ${\rm d}^{2}V/{\rm d}\phi^{2}$ dynamically turning negative, which was again found to be inconsequential for PBH production in models consistent with CMB data Suyama:2006sr . Finally, “axion” inflation coupled to gauge fields has also been revealed less prone to PBH production Caravano:2022epk than initially thought Linde:2012bt ; Garcia-Bellido:2016dkw .
@@
This was presented as a possibility already in Jedamzik:2010dq , but as far as we can infer, it was not based on an analysis of the non-linear evolution of the modes. We remind the reader that a well-accepted criterion for BH formation is that the hoop conjecture should be verified (see Ref. Misner:1973prb , pp. 867-868): the entirety of the object’s mass must be compressed to the point that it can fit within a sphere whose radius is equal to that object’s Schwarzschild radius. This is loosely equivalent to requiring that the Newtonian potential $|\Phi|\sim G{\cal M}/R$ associated to the mass ${\cal M}$ concentrated in a volume of radius $R$ attains $|\Phi|\sim 0.5$ (see e.g. Lyth:2005ze ). [^5] Some literature such as Cotner:2018vug ; del-Corral:2023apl has sometimes estimated PBH formation at sub-horizon scales relying on different criteria, which do not appear generically reliable. Notice that within a purely linear analysis, as the one in Jedamzik:2010dq , there is really no indication that the growth of density perturbations may lead to BH formation, since the curvature perturbation $\mathcal{R}$ (reducing in modulus to the gravitational potential $\Phi$ on sub-Hubble scales, see footnote 2) stays constant: Put otherwise, extrapolating the linear behaviour into the non-linear regime (a procedure which is not anyway to be trusted Parry:1998pn ) is not indicative of PBH formation.
In the appendices of Martin:2019nuw , the allegedly rather generic formation of PBHs was supported by solving the non-linear evolution equation of a spherically symmetric fluctuation of the energy density of the field, evolving in the homogeneous background of the scalar field itself, similarly to the approach followed in Goncalves:2000nz . As noted already in Goncalves:2000nz ; Martin:2019nuw , the solved equations are rather idealized, notably assuming spherically symmetric configurations. In the case of dust, the pioneering studies on the subject Khlopov:1980mg ; Polnarev:1985btg ; 1985MNRAS.215..575K showed that if and only if the fluctuations are homogeneous and spherically symmetric to a high-degree, the corresponding spherical symmetric solutions can be taken to represent reality.
@@
A number of articles have since been devoted to clarify the parametric dependence of the probability of the formation of (Hubble-sized) PBHs on the root-mean-square relative fluctuation of density on a mass scale ${\cal M}$, $\sigma({\cal M})$, when asphericity and angular momentum are taken into account Kuhnel:2016exn ; Harada:2016mhb ; Harada:2017fjm ; Carr:2018nkm ; Harada:2022xjp . In general, collapse is prevented by a number of effects: The isotropic approximation breaks down, gradients are important, non-Gaussian density and velocity effects are present, etc.
@@
In fact, the analogous problem in cold dark matter cosmology is well-known, and the dynamics (unless the power spectrum is extremely enhanced) leads to halo formation supported by velocity dispersion, rather than to BHs. The issue is then to understand to which extent something analogous happens also for the scalar field models of interest going beyond the dust-like fluid approximation, as noted e.g. in Martin:2020fgl . An outcome qualitatively analogous to the cold dark matter case has actually been recovered in numerical simulations of early structure formation, neglecting GR effects, in a universe subject to a quadratic inflaton potential, if the post-inflationary metastable state is assumed to survive long enough Musoke:2019ima . No sign of large gravitational potentials, which would cast doubts on resorting to a Newtonian approximation, is found there. Not surprisingly, the corresponding halos have a profile resembling an NFW one Eggemeier:2020zeg . Of course, since dealing with a bosonic degree of freedom, a solitonic condensate may form at the center of halos. Even in the rather idealised case where no self-interaction or other interactions are relevant, for these objects to collapse into BHs, the matter-dominated phase must last a very long time, equivalent to $\gtrsim 30\,$ e-folds Eggemeier:2021smj . Even neglecting the physical realism of these cases, definitely the PBH outcome is far from generic. [^6] Note that, as far as BH formation is concerned, the free bosonic case these simulations focus on is optimal, with self-interactions or interactions with other fields making formation of BH more challenging, see e.g. Garani:2021gvc .
@@
Is the GR treatment of gravity crucial to qualitatively alter these conclusions? Existing studies do not support this idea: For instance, the parametric family of potentials studied in Kou:2019bbc leads to BH formation only in a subset of cases when their parameter $\alpha<0.2$ (with $\alpha=1$ corresponding to the quadratic potential, see Fig. 8 in Kou:2019bbc ). The authors of Muia:2019coe also clearly state that the fate of non-linear structures forming in a quadratic inflaton potential is to disperse, not to end up forming BHs. This situation appears grim for PBH production, which are only found in special setups. In particular, we note that simulations leading to PBH formation start from already large density contrasts as initial conditions, whose dynamical plausibility in the post-inflationary era needs to be independently checked. We will return to this point in Sec. IV.